OpenStudy (blackops2_143):

What is the difference? Express your answer in lowest terms. -7/15 - (-11/15)= ?/?

OpenStudy (xapproachesinfinity):

hey

OpenStudy (blackops2_143):

hi

OpenStudy (xapproachesinfinity):

did you give it a try?

OpenStudy (wendy.ivette11714):

^

OpenStudy (xapproachesinfinity):

tell me where you stuck

OpenStudy (blackops2_143):

yes I have on every question but I get it wrong

OpenStudy (xapproachesinfinity):

ok no problem

OpenStudy (anonymous):

-7/15 - (-11/15)= -7/15 + 11/15

OpenStudy (wendy.ivette11714):

Can you show us where your stuck at?

OpenStudy (anonymous):

a double negative = 2 a positive

OpenStudy (anonymous):

so what do you have to do next?

OpenStudy (xapproachesinfinity):

eh please let the poster try understanding don't jump to answers

OpenStudy (anonymous):

ok sorry...

OpenStudy (xapproachesinfinity):

so dear poster what do you know about negative times negative

OpenStudy (xapproachesinfinity):

the problem is you are not helping if you just give away the answers :)

OpenStudy (blackops2_143):

not much, I moved to a new school and at my old one we weren't learning about it, kids here have been learning about this for a month and I just got here this week

OpenStudy (anonymous):

oh my

OpenStudy (wendy.ivette11714):

What have u done @blackops2_143 Can you show me the steps you took to solve this problem..

OpenStudy (xapproachesinfinity):

eh ok no problem

OpenStudy (wendy.ivette11714):

We can help you were you left off..

OpenStudy (anonymous):

well just tell us what is confusing u

OpenStudy (blackops2_143):

okay

OpenStudy (anonymous):

@blackops2_143 do u know how to work with fractions or negatives?

OpenStudy (blackops2_143):

fractions

OpenStudy (blackops2_143):

negatives are more confusing for me

OpenStudy (kendricklamar2014):

@Yuna123 Just let the user post the steps she took to solve this problem. Then we can all help her with this problem. Or else we wont get no where with this...

OpenStudy (anonymous):

ok @KendrickLamar2014

OpenStudy (anonymous):

sorry its just that i am in a hurry too so i will leave

OpenStudy (blackops2_143):

so I subtracted the fractions numerators but left the denominators and so now I don't know what to do

OpenStudy (wendy.ivette11714):

This is how I would solve this: \[\frac{ -7 }{ 15 } + \frac{ 11 }{ 15 }\] \[\frac{ -7 + 11 }{ 15 }\] \[-7 + 11 = ?\]

OpenStudy (kendricklamar2014):

^

OpenStudy (xapproachesinfinity):

its clear that when you multiply positive with positive you get positive 2x2=4 the issue is when multiplying two negatives -2x-2=4 why positive is the question consider 

OpenStudy (xapproachesinfinity):

that's one way of looking at product

OpenStudy (wendy.ivette11714):

What is: \[-7 + 11\] @blackops2_143

OpenStudy (blackops2_143):

so 4 ?

OpenStudy (xapproachesinfinity):

now if considered negative numbers

OpenStudy (wendy.ivette11714):

Yes, so the answer would be: \[\frac{ 4 }{ 15 }\]

OpenStudy (wendy.ivette11714):

Did you understand? @blackops2_143

OpenStudy (blackops2_143):

yes thank you both

OpenStudy (xapproachesinfinity):

i go 2 stpes left and 2 down i still got the same box that is of area 4 so -2x-2=4

OpenStudy (wendy.ivette11714):

Welcome =)

OpenStudy (xapproachesinfinity):

this explains why negative multiplied by negative is positive

OpenStudy (xapproachesinfinity):

so whenever you see -(-a) it is the same as -1x-a and that is equal to a a any positive number

OpenStudy (xapproachesinfinity):

i hope you got the picture

OpenStudy (xapproachesinfinity):

i m explain this since you said you have problem with negatives